OpenStudy (anonymous):

Find the exact value of the radical expression in simplest form. radical s^8+radical 25s^8+2radical s^8-radicals^4

OpenStudy (anonymous):

\[\sqrt{s^8}+\sqrt{25^8}+2\sqrt{s^8}+\sqrt{s^4}\]

OpenStudy (anonymous):

Help Please ! :(

OpenStudy (swissgirl):

ok lets start of with an example \( \sqrt{s^2} =? \)

OpenStudy (swissgirl):

\( = \sqrt{S*S} = S \) correct?

OpenStudy (swissgirl):

ok lets take more of a concrete example \( \sqrt{4} = \sqrt{2*2} = 2\)

OpenStudy (anonymous):

yea

OpenStudy (swissgirl):

So what wld \( \sqrt{s^4} =? \)

OpenStudy (swissgirl):

are you following?

OpenStudy (anonymous):

umm in confused lol sorry

OpenStudy (anonymous):

im*

OpenStudy (swissgirl):

hmmm so are you familiar with squareroots? like what is \( \sqrt{9} \)

OpenStudy (anonymous):

yes i am - 0

OpenStudy (swissgirl):

so what is \( \sqrt{9} =? \)

OpenStudy (anonymous):

is it 0?

OpenStudy (swissgirl):

no its not 0

OpenStudy (swissgirl):

http://www.mathsisfun.com/square-root.html

OpenStudy (swissgirl):

just read this short thing. It seriously clarifies everything

OpenStudy (anonymous):

okay thank you

OpenStudy (swissgirl):

but basically its the inverse of squaring \(3^2=9 \) so \( \sqrt{9} =3 \) The was i look at square root is there a number that when you multiply it by itself it will give you that original number

OpenStudy (swissgirl):

so for example \( \sqrt{16}\) What number when you multiply it by itself will give you 16

OpenStudy (swissgirl):

1*1=1 2*2=4 3*3=9 4*4=16

OpenStudy (swissgirl):

We see that when we multiply 4 buy itself it will give you 16 \( \sqrt{16}=\sqrt{4*4}=4

OpenStudy (swissgirl):

\( \sqrt{16} = \sqrt{4*4}=4 \)

OpenStudy (swissgirl):

ok Ill give you an example what is \( \sqrt{25}=?\)

OpenStudy (anonymous):

5

OpenStudy (swissgirl):

YAHOOOOOOO

OpenStudy (anonymous):

lol :)

OpenStudy (swissgirl):

hahah its hard to explain this topic when like its not in person

OpenStudy (swissgirl):

ok so lets start with \(\sqrt{s^4} \) how can that be rewritten?

OpenStudy (swissgirl):

\( \large \sqrt{s*s*s*s} = \sqrt{s^4} \) did you follow this?

OpenStudy (swissgirl):

@amistre64 maybe you can do a better job explaining

OpenStudy (amistre64):

i dunno if a can splain it better, but if you realize that radicals are frational exponents it might be better to see

OpenStudy (amistre64):

\[\large \sqrt[n]{X}=X^{\frac{1}{n}}\]

OpenStudy (amistre64):

\[\large \sqrt{s*s*s*s} = \sqrt{s^4}=s^{\frac{4}{2}}=s^2\]

OpenStudy (anonymous):

lol lol im sorry guys.

OpenStudy (swissgirl):

ohh this way is simplerrrrrrrr

OpenStudy (anonymous):

and yea i understood that!

OpenStudy (amistre64):

lets see hoe it plays then :) \[\sqrt{s^8}+\sqrt{25^8}+2\sqrt{s^8}+\sqrt{s^4}\] \[s^{8/2}+\sqrt{25}s^{8/2}+2s^{8/2}+s^{4/2}\] \[s^4+\sqrt{25}s^4+2s^4+s^2\]

OpenStudy (amistre64):

err... *how it plays out lol

OpenStudy (swissgirl):

you can pull out the 25

OpenStudy (amistre64):

you can, but i like to leave something to the imagination ;)

OpenStudy (swissgirl):

lol Thanks amistre

OpenStudy (amistre64):

enjoy ;)